OpenStudy (anonymous):

Which of the following is a point at which the tangent to y = -xe^(2x) is horizontal? The answer is (-0.5,0.18) where did the 0.18 come from?

OpenStudy (anonymous):

when u draw the graph u get y value 0.18

OpenStudy (zale101):

When you take the derivative of y = -xe^(2x) \(y'=-2xe^{2x}-e^{2x}\) you set it equal to zero \(y'=-2xe^{2x}-e^{2x}=\)\(-e^{2x}(2x+1)\) x= -0.5 Now, we have the x value of the tangent line, plug that back to the original equation

OpenStudy (zale101):

y = -xe^(2x) plug (-0.5) y=-(-0.5)e^(2(-0.5)) y=0.18...

OpenStudy (anonymous):

Oh! thank you so much, both of you

OpenStudy (anonymous):

try wolfram alpha to see the graph so u know for sure

OpenStudy (zale101):

You're welcome

OpenStudy (anonymous):

you will see something like this , it really helps when u can see it !